A car averages 12.5 L/100 km in city driving and 7.5 L/100 km in highway driving. In a week of mixed driving the car uses 35L of fuel and travelled 400 km. Determine the distance travelled in highway driving.
I'm having difficulty establishing the 2 equations. So far my attempts are
x + y = 400 12.5x + 7.5X = 8.75

Answer by scott8148(5896)

why two equations? __ let x="highway distance", so 400-x="city distance"

12.5L/100km=.125L/km __ 7.5L/100km=.075L/km

.075(x)+.125(400-x)=35 __ .075x+50-.125x=35 __ -.05x=-15